OpenStudy (madhu.mukherjee.946):

find the number of numbers less than 1000 and divisible by 5 which can be formed with the digits 0 1 2 3 4 5 6 7 8 9 without repetition

OpenStudy (madhu.mukherjee.946):

@imqwerty

OpenStudy (madhu.mukherjee.946):

@imqwerty fast :P

OpenStudy (madhu.mukherjee.946):

you are taking too long such a small sum:)

imqwerty (imqwerty):

a number divisible by 5 has either 5 at end or 0 at end so total such number will include all numbers of 1 digit , 2 digits... 1 digit nos are easy to count they are-5 2 digits nos - we consider 2 cases one with 0 at the end and one with 5 at the end A A] 0 at the end 1st place can be filled with-1,2,3,4,5,6,7,8,9 and 0 is fixed at 2nd place so total numbers=9x1=9 B]5 at end 1st place can be filled with-1,2,3,4,6,7,8,9=8nos we didn't take 0 cause then it wuld be a single digit no we also didn't take 5 cause its already set at 2nd place so total nos=8x1=8 total 2 digit nos=17 3 digit nos- A] 0 at end 1st place can be filled with-1,2,3,4,5,6,7,8,9=9nos 2nd place can be filled with 8 nos cause 1 number will be consumed in this^ step=8nos left to fill 2nd place 3rd place has 0=1no so total combinations =9x8 B] 5 at end 1st place can be filled with-1,2,3,4,6,7,8,9=8nos 2nd place can be filled with 8 nos cause we can take 0 here along with the 7 digits left after filling 1st place=8nos 3rd place has 5 =1no total nos=8x8 total 3 digit nos=72+64=136 total nos=136+17+9 =162

OpenStudy (madhu.mukherjee.946):

wrong answer is 154

imqwerty (imqwerty):

i wrote in brief :D

imqwerty (imqwerty):

lol wait

imqwerty (imqwerty):

hey i did that correct i jst made calculation mistake :P its 72+64+17+1=154

imqwerty (imqwerty):

i wrote that 9 by mistake :)P total i digit nos=1 total 2 digit nos=17 total 3 digit nos=72+64 total=154 >.<

imqwerty (imqwerty):

*1 digit nos

OpenStudy (madhu.mukherjee.946):

how did you get the 3 digit numbers pls explain

OpenStudy (ribhu):

the three digit numbers are obtained taking into account two cases

OpenStudy (madhu.mukherjee.946):

yeah 0 and 5

OpenStudy (madhu.mukherjee.946):

i got the 0 part just explain me the the 5 part

OpenStudy (ribhu):

first 0: the first digit can be filled with 9 choices.. second will be 8 and since 0 is fixed so we get 9*8 = 72 numbers ending in 0 with three digits

OpenStudy (madhu.mukherjee.946):

yeah even i got that

OpenStudy (ribhu):

see 5 is fixed.. so first digit can be filled in 8 ways..since 0 can not come in first place so second has 8 options..every number except for 5: so we have 8*8 = 64

OpenStudy (ribhu):